# Gravitational reduction of the wave function through the quantum theory of motion

### v1

@@ 3 The transition condition for a point particle
where, $\sigma_{0}$ is the initial width of the wave packet. By the short-time estimation, we mean $\sigma\approx\sigma_{0}$.
Now, we obtain the average of the equation (17) so that we can take the first step and find a criterion for the transition regime. We first derive the relations of the quantum potential and the quantum gravitational potential of the particle. The quantum potential is obtained as follows:
@@
The average values of the quantum force and the quantum gravitational forces are
@@
respectively. By equating these results, the characteristic width for which the transition from quantum to classical world occurs, is obtained as follows:
@@
By using the above relation, the critical mass is defined as
@@
This formula can be interpreted in such a way that for a specific width $\sigma_{0}$, if the particle mass is greater than the critical mass, the quantum gravitational force overcomes the quantum force. And if the particle mass is less than the critical mass, the quantum force overcomes the quantum gravitational force. And the particle behaves in a quantum way. That is, its wave packet spreads with time and never reduces. So, the uncertainty in the position of the particle increases. For example, consider a wave packet with the initial width $\sigma_{0}=10^{-2}cm$. The critical mass for this wave packet takes the value $m_{c}\approx 10^{-15}g$. In other words, for this wave packet, this value of mass can balance the quantum gravitational and quantum forces so that the wave packet does not spread. The different classifications of particle motion using the deviation equation (13) have been discussed in Ref.[33]. Nevertheless, by a combination of relations (24), (25) and (27), we derive a relation as follows:
@@
which classifies the possible regimes in this context.
@@
In the transition regime, where the averages of quantum force and quantum-gravitational force are balanced, we expect to have a stationary state for which the width of the wave packet is independent of time and the action or principal function, $S$, only depends on time. In other words, for a stationary state which can be represented in the form $\phi=\phi_{0}(\mathbf{x})e^{\frac{iS(t)}{\hbar}}$, the term $\nabla S(\mathbf{x},t)$ in the Hamilton-Jacobi equation (30) vanishes. So, we have
@@
Now, the minimizing of this equation i.e. $\delta\overline{E}=0$, gives $\sigma_{0}\approx\frac{\hbar^{2}}{Gm^{3}}$ which is the characteristic width for the transition regime. In this case, a stable system is formed, whose energy is given by
@@
where, it has been obtained through substituting the relation of the critical with into the relation (32). To obtain the critical values of an object too, we shall use the method of minimizing the Hamilton-Jacobi equation.
4 The transition condition for an object
For an object, we assume that the only degree of freedom is the coordinates of its center of mass. Thus, the associated wave packet is described by the same wave packet of a point particle. From fundamental physics, we know that for a homogeneous spherical mass distribution with radius $R$ and mass $m$, the classical gravitational potential is given by
@@ 4 The transition condition for an object
Now, we obtain $U_{\text{QG}}$ to estimate the transition condition for the object. The quantum gravitational energy, which is defined through the relation (14), becomes
@@
which we have taken the integral over the whole space, due to the nature of Gaussian distribution. In this problem, the average quantum potential energy of an object is the same as the average potential energy of a particle. So, the total energy of the object when we consider a stationary wave packet for the transition regime, is:
@@
Here, $\Sigma_{0}$ denotes the transition width for the object and $\sigma_{0}$ stands for the transition width of the point particle. This result agrees with the results of previous studies done in the framework of the standard quantum mechanics[13].
If we obtain the critical width through the equality of the averages of the quantum gravitational and quantum forces, we will be able to consider two limiting cases in this problem. The quantum gravitational force is as follows:
@@
This relation gives more reliable result when the quantum behavior is significant. In this situation, the wave packet is wide, and the object can be found outside the region $0\leq r\leq R$, most of the time.
For a macroscopic object, $\sigma_{0}\ll R$, and we get the following result,
@@
which is comparable to the relation of the gravitational force on a particle within a homogeneous mass distribution. For a macroscopic object, we expect the object to spend most of its time within the interval $0\leq r\leq R$.
For a macro object, we equate the average quantum gravitational force in relation (42) with the average of the quantum force to obtain the critical width of the wave packet. The result is the same as we saw in relation (38). For objects with the quantum-dominant behavior, for example a proton, the quantum dispersion is significant, and we use the relation (41) to obtain the critical width of the wave packet. This leads to
@@
This relation gives the value $\Sigma_{0}\approx 10^{6}\text{cm}$, for a proton, which is an acceptable value [8].
When, $\sigma_{0}=R$, the relations (38) or (43) give
@@
5 The reduction time for a point particle
We work in the gravity-dominant regime, where only quantum gravitational force is considered. The particle falls within its own quantum distribution $\varrho=\psi^{*}\psi=\mathcal{R}^{2}$, in configuration space. We can define the time needed for the particle to fall from an initial distance to the center of distribution. We shall see that the quantum motion of the particle is oscillatory in this regime, and the reduction time is proportional to the one-quarter of the period of oscillation in the short-time estimation. We shall also obtain the reduction time through the uncertainty in the particle’s quantum gravitational energy.
@@ 5 The reduction time for a point particle
$f_{\text{QG}}=-\sqrt{\frac{2}{\pi}}\frac{Gm^{2}}{\sigma^{3}_{0}}re^{-\frac{r^{2}}{2\sigma^{2}_{0}}}=m\ddot{r}.$ (46)
By assuming the initial conditions $\dot{r}(0)=0$ and $r(0)=\sigma_{0}$, and some approximations, it can be shown that the position of the particle obeys the equation
@@
The time needed for the particle from the initial distance $r(0)=\sigma_{0}$ to get $r(\tau)=0$, is proportional to the reduction time which is one-quarter of the period of motion [^1] In fact, it can be shown that the falling time is obtained through the relation $T_{\text{falling}}\approx\Big{(}\frac{\sigma_{0}^{3}}{Gm}\Big{)}^{\frac{1}{2}}F\Big{(}\frac{r_{f}}{r_{i}}\Big{)}$, where $F$ is the elliptical function. Here, $r_{i}$ and $r_{f}$ denote the initial and final distances respectively. For details, See Ref.[33]. According to the order of the numbers in this problem, we can simply state that the reduction time of the wave function is on the order of the oscillation period. According to the relation (49), we conclude that by increasing the mass of the particle for the fixed $\sigma_{0}$, the period of oscillation decreases and the particle gets the center of distribution faster. Also, the motion is harmonic, but it is not simple harmonic oscillation. Because, the period of motion depends on the initial position of the particle or the amplitude of the oscillation. It can be shown that the relation (49) is proportional to the reduction time of Karolyhazy in relation (2).
Equation (46) can be solved numerically to confirm the oscillatory motion of the particle. The results have been illustrated in Fig. (3) for some arbitrary values of involved parameters.
Figure 3: The oscillatory motion of the particle through its quantum distribution in configuration space with radius $\sigma_{0}$ and $r(0)=\sigma_{0}=1$, in a short-time estimation. The solid and dashed lines are related to the values $m=1$ and $m=30$, respectively. We have taken $\hbar=G=1$. The period of motion depends on the mass of the particle.
@@
Figure 4: The oscillatory motion of the particle through its quantum distribution in configuration space is not a simple harmonic motion. For both diagrams, we have taken $m=5$. But, for the dashed line, $r(0)=2$, while for the solid line, $r(0)=1$. We have taken $\hbar=G=1$. The period of motion depends on the amplitude of motion.
@@
$\overline{g_{\text{QG}}}\approx\frac{Gm}{\sigma_{0}^{2}},$ (51)
which is proportional to the average of the quantum gravitational acceleration which is obtained through quantum gravitational energy. According to these results and arguments, the time needed to get the particle to the center of the Gaussian distribution is proportional to the period of oscillation. So, we have
$\tau\approx\Big{(}\frac{\sigma_{0}^{3}}{Gm}\Big{)}^{\frac{1}{2}}=\Big{(}\frac{\sigma_{0}}{\overline{g_{\text{QG}}}}\Big{)}^{\frac{1}{2}},$ (52)
Since we have considered that $\sigma(t)\approx\sigma_{0}$ during the reduction time, we can substitute $\sigma_{0}=\frac{\hbar^{2}}{Gm^{3}}$ into relation (52) to get
@@
We can also estimate the reduction time through the uncertainty principle. When the particle falls into its own quantum distribution, from a distance $\sigma_{0}$, it reaches the center of distribution after time $\tau$. The average value of momentum at the time $\tau$, is $p=mu\approx mg\tau$. Now, according to $\Delta r\Delta p\sim\hbar$, we get $mg\tau\sigma_{0}\sim\hbar$. So, we will have
@@
One can see the whole story as follows. If, for a fixed probability distribution, the mass of the particle is less than the critical mass required to keep the probability distribution stationary, the particle will move away from the origin of the distribution under the effect of the quantum force and the corresponding wave function will never reduce. For a mass equal to the critical mass, the particle should form a stable system. Now, consider a fixed stationary quantum distribution. A particle with a mass greater than the critical mass falls within its quantum distribution. In fact, it has an oscillatory motion which is not simple. The collapse time is proportional to the period of oscillation. We can also look at the problem in the way that the particle has an uncertainty in its location. More uncertainty means more quantum behavior. If the mass of the particle exceeds the critical mass, the particle falls towards the center of the probability distribution and performs an oscillatory motion within its quantum probability distribution. In other words, the particle is trapped by gravity. And this confinement means that the location of the particle is determined with greater certainty. And as the mass increases, classical behavior also increases. This fluctuation also exists for the chair in our room, but it is so small that we know the location of the chair with high certainty. But the period of such an oscillation for a proton is so long that we have to wait thousands of years to observe the collapse of a proton.
Figure 5: The motion of the particle through its quantum distribution for several values of mass in the mixed regime. The initial condition are $r(0)=1$and $\dot{r}(0)=0$. We have also taken $\hbar=G=1$ and the initial width is $\sigma_{0}=1$.
We first solve this equation numerically. Then, we illustrate the particle motion in $(r,t)$ plane for different values of mass. Figure (5) shows that for a fixed initial width of the wave packet, by increasing the mass, the particle motion is oscillatory. While, by decreasing the mass, the quantum force becomes dominant and the particle moves away from the center of the probability distribution. So, we can not see the wave function reduction for light or elementary particles unless its wave function is reduced due to the measurement process. Of course, the measurement effect can also be related to the gravity caused by the measuring apparatus. Since the wave function of the measuring apparatus is entangled with the wave function of the quantum system, and the measuring apparatus is a macroscopic object, by reducing its wave function, the wave function of the quantum system will inevitably be reduced. This concept is well explained in Ref. [15].
6 The reduction time for an object
We realized that for a point particle, the reduction time is proportional to the period of oscillation with good approximation and associated relation (52), confirms the previous results which have been obtained in the framework of standard quantum mechanics. If we want to do the same for an object, we will find that the corresponding equation of motion is more complicated, and its solution cannot be obtained analytically unless we use some approximations. Therefore, we cannot easily obtain an analytical formula for the period of motion. However, a numerical method can be used to solve the equation of motion, and consequently, the period of oscillation can be obtained numerically. Nevertheless, we prefer to proceed the work through the uncertainty in the object’s quantum gravitational energy to estimate the reduction time of the wave function. To see the oscillatory motion of the object in the gravity-dominant regime, consider the equation of motion
@@ 6 The reduction time for an object
The solution of this equation for different values of parameters is seen in Fig. (6). The figure shows that the period of motion depends on the width of the wave packet. Also, the amplitude of motion is not the same above and below the horizontal axis. Because the first term in relation the (58) acts as a repulsive force. Naturally, this is due to the mass distribution and has nothing to do with the nature of the gravitational force. We did not have such a term for a point particle. Thus, for any arbitrary initial values, this equation can be solved and give us an estimation for the reduction time which is proportional to the period of oscillation.
Figure 6: The motion of the object within its quantum distribution in the gravity-dominant regime. We have assumed $R=1$. The mass of the object is $m=1$. We have also taken $\hbar=G=1$. For the dashed line, $\sigma_{0}=1.5$ and $r(0)=1.43$, while for the solid line, $\sigma_{0}=1$ and $r(0)=1.5$. It is obvious that the period of oscillation depends on the width of the wave packet.
According to these results, we conclude that the period of oscillation can be obtained numerically. One can calculate the time at which the motion diagram intersects the time axis to obtain the reduction time. Nevertheless, if we use the concept of uncertainty in energy, some approximated results can be obtained analytically. When, the object’s center of mass is at $r=\sigma_{0}$ in configuration space, an uncertainty of quantum gravitational energy is attributed to this distribution. We can determine the uncertainty in quantum gravitational energy during reduction time $\tau$, when the object’s center of mass has reached $r=0$. So, we use the relation (55) to get:
@@
where the values of $\alpha$ and $\beta$ depend on the order of approximation. For a ball with the size $R=5$ cm, the mass $m=100g$ and $\sigma_{0}\approx 10^{-12}cm$, the reduction time is of order of $10^{-23}$ s. The value of $\sigma_{0}$ has been estimated using relation (38). For a flea egg with a size of approximately $0.5mm$ and mass of about $10^{-5}g$, the reduction time is around $10^{-11}s$.
As relation (43) was obtained for a micro object by considering the asymptotic limit of the quantum gravitational force ($\sigma_{0}\gg R$), the asymptotic form of the quantum gravitational energy, should be considered here to provide a reliable result. For a micro object, the leading order of the asymptotic relation ($\sigma_{0}\gg R$) of the quantum gravitational energy is as follows:
@@
For a proton, this gives the value $\tau\sim 10^{15}$ s, which indicates that we do not see its collapse normally. This result is also in agreement with the results obtained in the framework of standard quantum mechanics. It seems the problem of gravitational reduction of the wave function in the framework of the Bohmian quantum mechanics can give reliable results. Therefore, this approach can also receive more attention.
7 Conclusions

### v2

@@ 2.1 The Transition Condition for a Point Particle
where, $\sigma_{0}$ is the initial width of the wave packet. By short-time estimation, we mean $\sigma\approx\sigma_{0}$. We shall see that the averages of gravitational force and gravitational potential depend on the width of the wave packet. Thus, in the short-time estimation, these quantities can be considered approximately constant. Now, we evaluate the average of equation (17) to find a criterion for the transition regime. We first derive the relations of the quantum potential and the quantum gravitational potential of the particle. The quantum potential is obtained as follows:
@@
respectively. The average values of these quantities are
@@
respectively. By equating these results, we obtain the characteristic width for which the transition from the quantum to the classical world occurs. The transition width is as follows:
@@
By using above relation, the critical mass is defined as follows:
@@
This formula can be interpreted such that for a specific width $\sigma_{0}$, if the particle mass is greater than the critical mass, the quantum gravitational force overcomes the quantum force. Conversely, if the particle mass is less than the critical mass, the quantum force overcomes the quantum gravitational force, and the particle behaves in a quantum manner. This means its wave packet spreads with time and never reduces, leading to an increase in the uncertainty of the particle’s position. For example, consider a wave packet with the initial width $\sigma_{0}=10^{-2}cm$. The critical mass for this wave packet is approximately $m_{c}\approx 10^{-15}g$. In other words, for this wave packet, this value of mass can balance the quantum gravitational and quantum forces so that the wave packet does not spread. By combination of relations (24), (25) and (27), we derive the following relation:
@@
This classification helps us understand the possible regimes in this context[34].
@@
In the transition regime, where the averages of quantum force and quantum-gravitational force are balanced, we expect to have a stationary state for which the width of the wave packet is independent of time and the action or principal function, $S$, only depends on time. In other words, for a stationary state that can be represented in the form $\psi=\psi_{0}(\mathbf{x})e^{\frac{iS(t)}{\hbar}}$, the term $\nabla S(\mathbf{x},t)$ in the Hamilton-Jacobi equation (30) vanishes. So, we have:
@@
Now, by minimizing this equation i.e., $\delta\overline{E}=0$, we obtain $\sigma_{0}\approx\frac{\hbar^{2}}{Gm^{3}}$, which is the characteristic width of the wave packet for the transition regime. In this case, the energy of the particle is given by
@@
which its negative sign indicates a bound system. To obtain the critical values of an object, we shall also use the method of minimizing the Hamilton-Jacobi equation.
2.2 The Transition Condition for an Object
For an object, we assume that the only degree of freedom is the coordinates of its center of mass. Thus, the associated wave packet is described by the same wave packet as a point particle. From fundamental physics, we know that for a homogeneous spherical mass distribution with radius $R$ and mass $m$, the classical gravitational potential is given by:
@@ 2.2 The Transition Condition for an Object
Now, we obtain $U_{\text{QG}}$ to estimate the transition width for the object. The quantum gravitational energy, which is defined through the relation (14), becomes:
@@
where we have taken the integral over the whole space due to the nature of the Gaussian distribution. Here, the average quantum potential of an object is the same as the average quantum potential of a particle. So, the average of total energy of the object when we consider a stationary wave packet for the transition regime is:
@@
Here, $\Sigma_{0}$ denotes the transition width for the object and $\sigma_{0}$ stands for the transition width of the point particle. This result agrees with the findings of previous studies conducted within the framework of standard quantum mechanics [13].
If we obtain the critical width through the equality of the averages of the quantum gravitational and quantum forces, we can consider two limiting cases in this problem. The quantum gravitational force is as follows:
@@
This relation gives a more reliable result when the quantum behavior is significant. In this situation, the wave packet is wide, and the object can be found outside the region $0\leq r\leq R$ most of the time.
For a macroscopic object, $\sigma_{0}\ll R$ and we get the following result,
@@
which is comparable to the relation of the gravitational force exerted on a particle within a homogeneous mass distribution. For a macroscopic object, we expect the object to spend most of its time within the interval $0\leq r\leq R$.
For a macroscopic object, we equate the average quantum gravitational force in relation (42) with the average quantum force to obtain the critical width of the wave packet. The result is the same as we saw in relation (38). For objects with quantum-dominant behavior, such as a proton, the quantum dispersion is significant. By equating the average quantum gravitational force in relation (41) with the average quantum force, we obtain the critical width of the wave packet. This leads to:
@@
This relation gives the transition width $\Sigma_{0}\approx 10^{6}\text{cm}$ for a proton, which is an acceptable value [8].
When $\sigma_{0}=R$, each of the relations (38) or (43) give
@@
In the next section, we discuss the meaning of the reduction time based on the quantum theory of motion.
3 The Reduction Time
We work in the gravity-dominant regime, where only quantum gravitational force is considered. The particle or object falls within its own quantum distribution $\varrho=\psi^{*}\psi=\mathcal{R}^{2}$, in configuration space. We can define the time needed for the particle to fall from an initial distance to the center of distribution. We shall see that the quantum motion of the particle or object is oscillatory in this regime, and the reduction time is proportional to one-quarter of the period of oscillation in the short-time estimation. The reduction time is also obtained through the uncertainty in the particle’s or object’s quantum gravitational energy.
3.1 The Reduction Time for a Point Particle
@@
$f_{\text{QG}}=-\sqrt{\frac{2}{\pi}}\frac{Gm^{2}}{\sigma^{3}_{0}}re^{-\frac{r^{2}}{2\sigma^{2}_{0}}}=m\ddot{r},$ (46)
where, we have used the relation (23). By assuming the initial conditions $\dot{r}(0)=0$ and $r(0)=\sigma_{0}$, and some approximations, it can be shown that the position of the particle obeys the equation
@@
The time needed for the particle to travel from the initial distance $r(0)=\sigma_{0}$ to $r(\tau)=0$ is proportional to the reduction time, which is one-quarter of the period of motion [^1] In fact, it can be shown that the falling time is obtained through the relation $T_{\text{falling}}\approx\Big{(}\frac{\sigma_{0}^{3}}{Gm}\Big{)}^{\frac{1}{2}}F\Big{(}\frac{r_{f}}{r_{i}}\Big{)}$, where $F$ is the elliptical function. Here, $r_{i}$ and $r_{f}$ denote the initial and final distances respectively. For details, See Ref.[33]. It can be shown that relation (49) is proportional to the reduction time in relation (2). We can simply state that the reduction time of the wave function is proportional to the period of oscillation. According to relation (49), we conclude that by increasing the mass of the particle for the fixed $\sigma_{0}$, the period of oscillation decreases and the particle reaches the center of distribution faster. Additionally, the motion is harmonic, but it is not simple harmonic oscillation because the period of motion depends on the initial position of the particle or the amplitude of the oscillation.
Equation (46) can be solved numerically to confirm the oscillatory motion of the particle. The results are illustrated in Fig. (2) for some arbitrary values of the involved parameters. The reduction of the wave function does not mean that the particle remains at one point after the collapse of its wave function. Rather, the particle oscillates within an interval that is inversely related to its mass. This interval is so small for a classical particle that we confidently say the particle has a definite location. On the other hand, for a microscopic particle, this interval can be such that the oscillation period lasts for thousands of years, and the uncertainty in the location of the particle is significant.
Figure 2: The oscillatory motion of the particle through its quantum distribution in configuration space with initial condition $r(0)=\sigma_{0}=1$, in a short-time estimation. The solid and dashed lines correspond to the values $m=1$ and $m=30$, respectively. Parameters are set to $\hbar=G=1$. The period of motion depends on the mass of the particle.
@@
Figure 3: The oscillatory motion of the particle through its quantum distribution in configuration space is not a simple harmonic motion. For both diagrams, $m=5$. For the dashed line, $r(0)=2$, while for the solid line, $r(0)=1$. Parameters are set to $\hbar=G=1$. The period of motion depends on the amplitude of motion.
@@
$\overline{g_{\text{QG}}}\approx\frac{Gm}{\sigma_{0}^{2}}.$ (51)
According to these results and arguments, we have:
$\tau\approx\Big{(}\frac{\sigma_{0}^{3}}{Gm}\Big{)}^{\frac{1}{2}}=\Big{(}\frac{\sigma_{0}}{\overline{g_{\text{QG}}}}\Big{)}^{\frac{1}{2}}.$ (52)
Since we have considered that $\sigma(t)\approx\sigma_{0}$ during the reduction time, we can substitute $\sigma_{0}=\frac{\hbar^{2}}{Gm^{3}}$ into relation (52) to obtain:
@@
Figure 4: Motion of the particle through its quantum distribution for various mass values in the mixed regime. Initial condition are $r(0)=1$and $\dot{r}(0)=0$, with $\hbar=1$, $G=1$, and initial width $\sigma_{0}=1$.
Figure (4) shows that for a fixed initial width of the wave packet, by increasing the mass, the particle motion becomes oscillatory. However, by decreasing the mass, the quantum force becomes dominant and the particle moves away from the center of the probability distribution. Thus, we cannot observe the wave function reduction for light or elementary particles unless their wave function is reduced due to the measurement process. The measurement effect can also be related to the gravity caused by the measuring apparatus. Since the wave function of the measuring apparatus is entangled with the wave function of the quantum system, and the measuring apparatus is a macroscopic object, by reducing its wave function, the wave function of the quantum system will inevitably be reduced. This concept is well explained in Ref. [15].
One can see the whole story as follows:
If, for a fixed probability distribution, the mass of the particle is less than the critical mass required to keep the probability distribution stationary, the particle will move away from the origin of the distribution under the effect of the quantum force and the corresponding wave function will never reduce. For a mass equal to the critical mass, the particle should form a stable system. Now, consider a fixed stationary quantum distribution. A particle with a mass greater than the critical mass falls within its quantum distribution. In fact, it has an oscillatory motion which is not simple. The collapse time is proportional to the period of oscillation. We can also look at the problem in the way that the particle has an uncertainty in its location. More uncertainty means more quantum behavior. If the mass of the particle exceeds the critical mass, the particle falls towards the center of the probability distribution and performs an oscillatory motion within its quantum probability distribution. In other words, the particle is trapped by gravity. This confinement means that the location of the particle is determined with greater certainty. As the mass increases, classical behavior also increases. This fluctuation also exists for the chair in our room, but it is so small that we know the location of the chair with high certainty. However, the period of such an oscillation for a proton is so long that we have to wait thousands of years to observe the collapse of a proton.
We can also estimate the reduction time through the uncertainty principle. When the particle falls into its own quantum distribution from a distance $\sigma_{0}$, it reaches the center of distribution after time $\tau$. The average value of momentum at the time $\tau$, is $p=mu\approx mg\tau$. Now, according to $\Delta r\Delta p\sim\hbar$, we get $mg\tau\sigma_{0}\sim\hbar$. So, we will have
@@
3.2 The Reduction Time for an Object
We realized that for a point particle, the reduction time is proportional to the period of oscillation with good approximation, and the associated relation (52) confirms the previous results obtained in the framework of standard quantum mechanics. If we want to do the same for an object, we find that the corresponding equation of motion is more complicated, and its solution cannot be obtained analytically unless we use some approximations. Therefore, we cannot easily derive an analytical formula for the period of motion. However, the equation of motion can be solved numerically, and the period of oscillation can be evaluated. To observe the oscillatory motion of the object in the gravity-dominant regime, consider the equation of motion as follows:
@@ 3.2 The Reduction Time for an Object
The solution of this equation for different values of parameters is seen in Fig. (5). The figure shows that the period of motion depends on the initial position of the object. Also, the amplitude of motion is not the same above and below the horizontal axis because the first term in the parentheses of relation (58) acts as a repulsive force.
Figure 5: The motion of the object within its quantum distribution in the gravity-dominant regime. Parameters are set to $R=1$, $m=1$, $\hbar=1$, and $G=1$. For the dashed line, $\sigma_{0}=1.5$ and $r(0)=1.43$; for the solid line, $\sigma_{0}=1$ and $r(0)=1.5$. The period of oscillation is dependent on the initial position of the object and the width of its associated wave packet.
Naturally, this is due to the mass distribution and has nothing to do with the nature of the gravitational force. We did not have such a term for a point particle. For various initial conditions, equation (58) can be solved to give us an estimation for the reduction time, which is proportional to the period of oscillation.
While solving the above equation provides the desired solutions for objects of varying sizes and masses and allows us to estimate the reduction time, it is intriguing that by considering the uncertainty in the particle’s quantum gravitational energy, we also derive an approximated analytical relation for the reduction time. When, the object’s center of mass is at $r=\sigma_{0}$ in configuration space, an uncertainty in quantum gravitational energy can be attributed to this distribution. In fact we consider the uncertainty in quantum gravitational energy during the reduction time $\tau$, when the object’s center of mass has reached $r=0$. So, we use the relation (56) to get:
@@
where the values of $\alpha$ and $\beta$ depend on the degrees of approximation. For a ball with the size $R=5$ cm, the mass $m=100g$ and $\sigma_{0}\approx 10^{-12}cm$, the reduction time is on the order of $10^{-23}s$. The value of $\sigma_{0}$ has been estimated using relation (38). In fact, we estimate the critical width $\Sigma_{0}$ for an object. For a flea egg with a size of approximately $0.5mm$ and a mass of about $10^{-5}g$, the reduction time is around $10^{-11}s$.
As relation (43) was obtained for a micro object by considering the asymptotic limit of the quantum gravitational force ($\sigma_{0}\gg R$), the asymptotic form of the quantum gravitational energy should be considered here to provide a reliable result for the reduction time. For a micro object, the leading order of the asymptotic relation ($\sigma_{0}\gg R$) of the quantum gravitational energy is as follows:
@@
For a proton, this gives the value $\tau\sim 10^{15}s$ , which indicates that we do not see its collapse normally. This result is also in agreement with the results obtained in the framework of standard quantum mechanics. In Bohmian quantum mechanics, the position of a particle or object is a hidden variable, meaning we do not know which trajectory is chosen during its motion. However, when the particle moves along any trajectory within the ensemble, it must obey the equation (7). Assuming the initial position of the particle or object is near the initial wave packet width, we find that the approximated analytical results align with those obtained from standard quantum mechanics. However, the deterministic structure of Bohmian quantum mechanics, which accounts for underlying processes, allows us to solve the problem for any initial position and velocity derived from the initial wave packet.
Figure 6: The motion of the object within its quantum distribution in the gravity-dominant regime for different values of mass. Parameters are set to $R=1$, $\sigma_{0}=1$, $\hbar=1$, and $G=1$. The initial position and velocity are $r(0)=0.1$ and $r^{\prime}(0)=-0.1$ respectively. For the dashed line, $m=2$; for the solid line, $m=4$. The period of oscillation is inversely dependent on the mass of the object.
In Fig. 6, we assume the object has an initial velocity towards the center of the distribution and examine how the period of motion changes with variations in mass, while keeping the wave packet width and the size of the object constant. As expected, increasing the mass decreases the period of motion, allowing the object to reach the center of the distribution more quickly.
According to these investigations and results, the problem of gravitational reduction of the wave function within the framework of Bohmian quantum mechanics appears to yield reliable results. Consequently, this approach merits further attention.
4 Conclusion